OpenStudy (anonymous):

The difference of two positive numbers is 8. The larger is twice the smaller decreased by 3. Find the two numbers. @eashy

OpenStudy (anonymous):

let x be the larger number, y be the smaller number x-y=8 x=2y-3

OpenStudy (anonymous):

use the substitution method to find x and y

OpenStudy (anonymous):

2y-3-y=8

OpenStudy (anonymous):

looks good :) now add 3 to both sides so that the y's are on the left and the numbers on the right

OpenStudy (anonymous):

2y-y=11

OpenStudy (anonymous):

now simply the left side to find y then substitute the value you found for y into x=2y-3 to get x

OpenStudy (anonymous):

y=11! lol x=2(11)-3 x=22-3 x=19!

OpenStudy (anonymous):

excellent. the numbers are 11 and 19

OpenStudy (anonymous):

Thanks!